Five hundred eleven (511) homes in a certain southern California community are randomly surveyed to determine if they meet minimal earthquake preparedness recommendations. One hundred seventy-three (173) of the homes surveyed met the minimum recommendations for earthquake preparedness and 338 did not.

The point estimate for the population proportion of homes that do not meet the minimum recommendations for earthquake preparedness is _____________.

(Write as a decimal and round to the 4th decimal place)
Answer #1

\hat{p}=X/n

\hat{p}=338/571=0.66

#The point estimate for the population proportion of homes that do not meet the minimum recommendations for earthquake preparedness is 0.66